OpenStudy (anonymous):

Write a formula for the general term (the nth term) of the geometric sequence. 1/2 ; -(1/10) ; 1/50 ; -(1/250), .....

OpenStudy (anonymous):

A. an = (1/2)n - 1(3/5 B. an = 1/2 - 1/5(n - 1) C. an = 1/2(-1/5)n - 1 D. an = 1/5(-1/2)n - 1